OpenStudy (anonymous):

A pyramid with a square base that is 10 cm on a side has a height of 12 cm. The volume of the pyramid is______ cm3.

OpenStudy (aroub):

V= 1/3 bh V= 1/3 (10)(12) V= 40 cm^3